OpenStudy (anonymous):

Need The Full Steps On How To Combine Like Terms and Simplify This Problem: 4/5(2z+10)-1/2(z+3)

OpenStudy (anonymous):

multiply it out to obtain 8/5z + 8 -1/2z - 3/2 then add terms with z and constants together

OpenStudy (anonymous):

how do i show the work to get 8/5z+8-1/2z-3/2

OpenStudy (campbell_st):

is the question 1. \[\frac{4}{5} (2z + 10) - \frac{1}{2}(z + 3)\] or is it 2. \[\frac{4}{5(2z + 10)} - \frac{1}{2(z + 3)}\]

OpenStudy (anonymous):

the 1st one campbell

OpenStudy (anonymous):

4/5 * 2z = 8/5z ...etc

OpenStudy (campbell_st):

ok... so you are adding and subtracting fractions \[( \frac{8}{5} - \frac{1}{2})z + (8 - \frac{3}{2}) \] now just simplify for the answer.

OpenStudy (anonymous):

Campbell how do i show work to show that i got 8/5

OpenStudy (campbell_st):

well it \[\frac{4}{5} \times 2 = \frac{8}{5} \] you are distributing

OpenStudy (anonymous):

Now whaT

OpenStudy (anonymous):

Campbell

OpenStudy (campbell_st):

ok... so the same for the 10 \[\frac{4}{5} \times 10 = 8\] next look at the 2nd set of brackets \[-\frac{1}{2} \times 1 = \frac{1}{2} z\] and \[- \frac{1}{2} \times 3 =- \frac{3}{2}\]

OpenStudy (anonymous):

I'm So Confused I Suck At Math Like Big Time

OpenStudy (campbell_st):

ok... here are the bits again distributing means multiplying the outside by everything inside. \[\frac{4}{5} \times 2z = \frac{8}{5} z\] \[\frac{4}{5} \times 10 = 8\] \[-\frac{1}{2} \times z =- \frac{1}{2} z\] \[- \frac{1}{2} \times 3 = \frac{3}{2}\] so after distributing you get \[\frac{8}{5} z + 8 - \frac{1}{2} z - \frac{3}{2} \] like terms have the same letter and can be added or subtracted then you will get \[(\frac{8}{5} - \frac{1}{2})z + ( 8 - \frac{3}{2}) \] put the fractions over the same denominators will give \[(\frac{16}{10} - \frac{5}{10}) z + ( \frac{16}{2} - \frac{3}{2})\] this simplifes to \[\frac{11}{10} z + \frac{13}{2} \] change the improper fractions to mixed numbers.. \[1 \frac{1}{10} z + 6 \frac{1}{2} \]